OpenStudy (anonymous):

Help me i will award medal and fan! The population of a local species of dragonfly can be found using an infinite geometric series where a1 = 42 and the common ratio is three fourths. Write the sum in sigma notation and calculate the sum that will be the upper limit of this population.

OpenStudy (anonymous):

the summation of three fourths times 42 to the i minus 1 power, from i equals 1 to infinity.; the sum is divergent. the summation of three fourths times 42 to the i minus 1 power, from i equals 1 to infinity.; the sum is 168. the summation of negative 42 times three fourths to the i minus 1 power, from i equals 1 to infinity.; the sum is divergent. the summation of negative 42 times three fourths to the i minus 1 power, from i equals 1 to infinity.; the sum is 168.

OpenStudy (anonymous):

Those are my answer choices

OpenStudy (kirbykirby):

actually hold on... -_-

OpenStudy (anonymous):

haha okay and it wants to know if its divergent or the sum is 168

OpenStudy (anonymous):

cant believe i have kirbykirby working on my problme xD

OpenStudy (kirbykirby):

Well it will be convergent for sure since the ratio is a fraction less than 1

OpenStudy (anonymous):

oh true so that removes two of the answer choices

OpenStudy (kirbykirby):

Actually, before I continue.. does your prof write the geometric series starting from \(a_1\) or \(a_0\)? Like do they write \(a_0+a_1+a_2+...\) or \(a_1+a_2+...\) This will make a difference

OpenStudy (anonymous):

No i wouldnt think so since i am in highschool, so im pretty sure they dont do that

OpenStudy (kirbykirby):

?

OpenStudy (kirbykirby):

Which one do they use

OpenStudy (anonymous):

usually a1

OpenStudy (anonymous):

i have never seen a0

OpenStudy (kirbykirby):

ok

OpenStudy (kirbykirby):

The geometric series is written as: \[\large 42+42\left( \frac{3}{4}\right)+42\left( \frac{3}{4}\right)^2+42\left( \frac{3}{4}\right)^3+...\\ =\large a_1+a_1r+a_1r^2+a_1r^3+...\\ =\large a_1+a_2+a_3+a_4...\] Now, the geometric series in sigma notation is written as: \[ \large\sum_{i=0}^{\infty}a_1r^i\] \[\large \sum_{i=0}^{\infty}42\left(\frac{3}{4} \right)^{i}=\sum_{i=1}^{\infty}42\left( \frac{3}{4}\right)^{i-1} =\large \frac{42}{1-\frac{3}{4}}\]

OpenStudy (anonymous):

Thank you so much this was so helpful, you earned yourself a fan and a medal

OpenStudy (kirbykirby):

:) np